OpenStudy (anonymous):

what is the slope of the line that is perpendicular to the line that passes through the points (7,6) and (3,-4)

OpenStudy (anonymous):

ok so what u do is:

OpenStudy (lncognlto):

m = y2-y1/x2-x1 = (-4-6)/(3-7) =-10/-4 =5/2 therefore, as m(perp) x m = -1, m(perp) = -2/5

OpenStudy (anonymous):

Thank you so much!!!!!!!

OpenStudy (lncognlto):

You're welcome! :)

OpenStudy (anonymous):

perpendicular slopes have negative reciprocals- remember that

OpenStudy (anonymous):

Ok I'll try

OpenStudy (anonymous):

:)